OpenStudy (anonymous):

Solve. 3x2 = 33x – 27

OpenStudy (anonymous):

\[3x^2-33x+27=0\] the roots are: \[[-(-33)\pm \sqrt{33^2-4*3*27}]/2*3\]

OpenStudy (anonymous):

\[x1=(33+\sqrt{765})/6\] \[x2=(33-\sqrt{765})/6\]

OpenStudy (anonymous):

what is the answer?

OpenStudy (anonymous):

because for my answer choices i cant seem to figure which is correct

OpenStudy (anonymous):

?